12. Provide an appropriate response. Samples of size n- 240 are randomly selected from the population of numbers (0 through 20) produced by a random-number generator, and the variance is found for each sample. What is the distribution of the sample variances? O normal (approximately) O skewed to the right O skewed to the left O not enough information provided 13. Choose the correct response. ( point) Why is sampling without replacement acceptable with a large population? When a small sample is taken from a large population, the samples maintain their independence O When a large sample is take from a large population, the sample retains the characteristics of the original distribution When a sample without replacement is taken, there is no requirement to maintain the same sample size When a sample without replacement is taken, the samples will have the appropriate amount of data
0. Provide an appropriate response. (1 point) Find the area of the shaded region. The graph depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 (as on the Wechsler test). 40 50 0 70 80 100 110 120 130 140 150 160 170 0.4400 O 0.6293 0.8051 0.7486 11. Select the correct response (1 point) What are unbiased estimators targeting? sample means independent samples population parameters
8. Find the indicated probability. (l point) In one region, the September energy consumption levels for single-family homes are found to be normally distributed with a mean of 1050 kWh and a standard deviation of 218 kWh. For a randomly selected home, find the probability that the September energy consumption level is between 1100 kWh and 1225 kWh О 0.1971 0.3791 0.2881 0.0910 9. Provide an appropriate response. (I point Find the area of the shaded region. The graph depicts IQ scores of adults, and those scores are normally distributed with a mean of 100 and a standard deviation of 15 (as on the Wechsler test). 6s 70 73 eo05 90 95 100 105110115 120 125130135 140 i45 0.7303 0.7619 0.7938 0.7745
5. If z is a standard normal variable, find the probability. (1 point) The probability that z lies between 0 and 3.01 00.9987 О 0.1217 0 0.5013 0 0.4987 6. Solve the problem. Round to the nearest tenth unless indicated otherwise. (1 poirt) In one region, the September energy consumption levels for single-family homes are found to be normally distributed with a mean of 1050 kWh and a standard deviation of 218 kWh. Find P45, which is the consumption level separating the bottom 45% from the top 55% 01021.7 1078.3 О 1087.8 1148.1
Answer #1

#5 By using standard normal table (shown below), the probability that z lies between 0 and 3.01 is:

P(0 < z < 3.01) = P(z < 3.01)-P(z < 0) = 0.9987-0.5000-0.4987

Standard Normal Cumulative Probability Table 0.00 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 0.0 0.5000 0.5040 0.5080 0.5120 0.5160 0.5199 0.5239 0.52790.5319 0.5359 0.53980.5438 0.5478 0.5517 0.5557 0.55960.5636 0.5675 0.5714 0.5753 0.5793 0.5832 0.5871 0.5910 0.59480.5987 0.6026 0.60640.6103 0.6141 0.6179 0.6217 0.6255 0.6293 0.6331 0.6368 0.6406 0.6443 0.6480 0.6517 0.65540.6591 0.6628 0.66640.6700 0.6736 0.6772 0.6808 0.6844 0.6879 0.1 0.2 0.3 0.4 3.0 3.2 3.3 3.4 0.9987 10.9987 0.9987 0.9988 0.9988 0.9989 0.9989 0.9989 0.9990 0.9990 0.9990 0.99910.9991 0.9991 0.99920.9992 0.9992 0.9992 0.99930.9993 0.9993 0.99930.99940.99940.99940.99940.99940.99950.9995 0.9995 0.9995 0.9995 0.9995 0.9996 0.9996 0.9996 0.99960.9996 0.99960.9997 0.9997 0.99970.9997 0.9997 0.99970.99970.99970.9997 0.99970.9998

--------------------------------------------------------------------------------------------

#6 Look into standard normal table (shown below) for area closest to 0.4500, which is 0.4483. The z score corresponding to this is -0.13. Therefore, we can say, z score corresponding to P45 is -0.13.

z=rac{x-mu}{sigma}Rightarrow x=mu+zsigma =1050-0.13(218)=1021.66approx {color{Red} 1021.7}

Standard Normal Cumulative Probability Table Cumulative probabilities for NEGATIVE z-values are shown in the following table: 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.00 0.08 0.09 -0.4 0.3446 0.3409 0.3372 0.3336 0.3300 0.3264 0.3228 0.3192 0.3156 0.3121 -0.3 0.3821 0.3783 0.3745 0.3707 0.3669 0.3632 0.35940.3557 0.3520 0.3483 -0.2 0.4207 0.4168 0.4129 0.4090 0.4052 0.4013 0.3974 0.3936 0.3897 0.3859 0.4602 0.4562 0.4522 0.44830.4443 0.4404 0.43640.4325 0.42860.4247 0.0 0.5000 0.4960 0.4920 0.4880 0.4840 0.4801 0.4761 0.4721 0.46810.4641 -0.1

--------------------------------------------------------------------------------------------

#8 By using standard normal table (shown below), the probability that September energy consumption level is between 1100 kWh and 1225 kWh is:

Pl1100 < r < 1225)-P(11002181050 r-μ 1222 P(11001225)-P 218 = P(0.23 < z < 0.80) P(z < 0.80)-P(z < 0.23) 0.7881-0.5910 0.1971

Standard Normal Cumulative Probability Table Cumulative probabilities for POSITIVE z-values are shown in the following table: 0.01 0.09 0.00 0.5000 0.5040 0.5080 0.5120 0.5160 0.51990.5239 0.5279 0.5319 0.5359 0.53980.5438 0.5478 0.5517 0.5557 0.55960.5636 0.5675 0.5714 0.5753 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.0 0.1 0.2 | 0.5793 0.5832 0.5871 |0.5910| 0.5948 0.5987 0.6026 0.6064 0.6103 0.6141 0.3 0.4 0.5 0.6 0.7 0.8 0.78810.7910 0.79390.7967 0.7995 0.8023 0.8051 0.8078 0.81060.8133 0.9 0.6179 0.6217 0.6255 0.6293 0.6331 0.6368 0.6406 0.6443 0.6480 0.6517 0.65540.6591 0.6628 0.66640.6700 0.6736 0.6772 0.6808 0.6844 0.6879 0.6915 0.6950 0.6985 0.70190.70540.7088 0.7123 0.7157 0.7190 0.7224 0.7257 0.7291 0.7324 0.7357 0.73890.7422 0.74540.7486 0.7517 0.7549 0.7580 0.7611 0.7642 0.7673 0.77040.77340.7764 0.77940.7823 0.7852 0.8159 0.8186 0.8212 0.8238 0.8264 0.8289 0.8315 0.8340 0.8365 0.8389

--------------------------------------------------------------------------------------------

#9 By using standard normal table (shown below), the probability that IQ score is between 85 and 125 is:

risS <バ125)-P(8515100 < r-μ < 12-100) P(85 < r < 125)- P (-15 = P(-1.00 < z < 1.67) P(z < 1.67)-P(z <-1.00) 0.9525-0.1587-0.7

Standard Normal Cumulative Probability Table 0.00 0.0808 0.0793 0.0778 0.07640.0749 0.0735 0.0721 0.0708 0.0694 0.0681 0.0968 0.09510.0934 0.0918 0.0901 0.0885 0.08690.0853 0.0838 0.0823 0.1151 0.1131 0.1112 0.1093 0.1075 0.1056 0.1038 0.1020 0.1003 0.0985 0.1357 0.1335 0.1314 0.1292 0.1271 0.125 0.1230 0.1210 0.1190 0.1170 0.01 0.02 0.03 0.04 0.05 0.06 0.07 0.08 0.09 1.3 -1,0 | О. 87 0.1562 0.1539 0.1515 0.1492 0.1469 0.1446 0.1423 0.1401 0.1379. 1.5 1.6 1.7 1.8 1.9 0.9332 0.9345 0.93570.9370 0.9382 0.93940.9406 0.9418 0.9429 0.9441 0.9452 0.9463 0.9474 0.9484 0.9495 0.9505 0.9515 0.9525 0.9535 0.9545 0.95540.95640.9573 0.9582 0.95910.9599 0.9608 0.9616 0.9625 0.9633 0.96410.9649 0.9656 0.9664 0.9671 0.9678 0.9686 0.9693 0.96990.9706 0.9713 0.9719 0.9726 0.9732 0.9738 0.9744 0.9750 0.9756 0.97610.9767

--------------------------------------------------------------------------------------------

#12 Since variance cannot be negative, the sample variance values must be 0 or above. Therefore, the distribution of sample variances must be skewed to the right .

--------------------------------------------------------------------------------------------

#13 When small sample (less than 5% of the population) is taken from the population, the samples maintain their independence. Therefore, correct answer is first option .